Why does the graph of a quadratic intercept the y-axis at only one point? Does the graph of a quadratic always intercept the x-axis in two points? Give examples.


Answer by venugopalramana(3286)

Why does the graph of a quadratic intercept the y-axis at only one point? Does the graph of a quadratic always intercept the x-axis in two points? Give examples.
NO IT DEPENDS ON THE EQN.IT MAY CUT AT 2 OR 1 OR NO POINT
EX.1.....X^2-5X+6=0=(X-2)(X-3)=0
IT CUTS X AXIS AT X=2 AND X=3
EX.2.....X^2-4X+4=0=(X-2)^2=0
IT CUTS X AXIS AT ONE POINT ONLY AT X=2
EX.3....X^2+4=0
IT DOES NOT CUT THE X AXIS AT ALL